OpenStudy (anonymous):

Suppose you are given two quadrilaterals, ABCD~ EFGH. If the similarity ratio of ABCD to EFGH is 5:3 and EH = 60, find AD.

Directrix (directrix):

Lengths of corresponding sides of similar triangles are in proportion. ABCD~ EFGH means that AD is to EH as 5 is to 3 because 5:3 is the ratio of corresponding sides. AD/EH = 5/3 EH = 60 So, the proportion becomes AD/60 = 5/3. Cross multiply and solve for AD. Post what you get. @phillitup

OpenStudy (anonymous):

100?

Directrix (directrix):

That is what I got.

OpenStudy (anonymous):

thanks!

Directrix (directrix):

You're welcome.